By 3:00 PM, John had plowed 7 driveways and by 5:30 PM he had completed 13.
a) Find John’s plowing rate in number of driveways per hour.
b) Find John’s plowing rate in minutes per driveway

Answer by stanbon(75887)

By 3:00 PM, John had plowed 7 driveways and by 5:30 PM he had completed 13.
a) Find John’s plowing rate in number of driveways per hour.
---
rate = (# of driveways)/(# of hours) = 6/2.5 = 2.4 driveways/hour
----------------------------------------------------------
b) Find John’s plowing rate in minutes per driveway
(2.4)(driveways/hour)(1 hr/60 minutes)
---
= (2.4*1)/(1*60)(driveways/minute)
---
= 0.04 driveways/minute
============================
cheers,
Stan H.